Find the molar mass of an unknown gas which diffuses twice as fast as Cl2 gas.
1
Expert's answer
2020-12-01T14:26:10-0500

Molar mass of Cl2, M1 = 35.5x2=71.0gmol-1

Let the molar mass of the unknown gas X be M2

from Graham's law of diffusion,

Rx/Rcl = √M1/M2

2/1=√71/M2

2²/1²= 71/M2

4M2= 71

M2 = 71/4= 17.74gmol-1